OpenStudy (anonymous):

does f(9) exist at (x-9)^2/3

OpenStudy (jdoe0001):

well.. what does it give you?

OpenStudy (jdoe0001):

just plug it in

OpenStudy (anonymous):

0 so no?

OpenStudy (jdoe0001):

... well. 0 then :) so it exists, is 0 \(\bf \large (x-9)^{\frac{2}{3}}\implies \sqrt[3]{(x-9)^2}\implies \sqrt[3]{(0)^2}\implies 0\)

OpenStudy (anonymous):

i'm confused how does it exist @jdoe0001 ?

OpenStudy (jdoe0001):

\(\bf \large \sqrt[3]{0^2}\implies \sqrt[3]{0}\implies \sqrt[3]{0^3}\implies 0\)

OpenStudy (jdoe0001):

is not undefined, thus it exists, it may not be a nice looking integer, it's just 0

OpenStudy (anonymous):

oh ok would you mind helping me with somehing else @jdoe0001 ?

OpenStudy (jdoe0001):

well, you can always post anew, so we can all see it and help and revise each other

OpenStudy (anonymous):

ok